OpenStudy (anonymous):

A sample of chlorine gas has a pressure of 7.25 kPa at 20.0°C. What will its pressure be at 60.0°C if its volume remains constant? A. 2.42 kPa B. 8.24 kPa C. 21.8 kPa D. 6.38 kPa

OpenStudy (anonymous):

@JFraser i think its (C)

OpenStudy (jfraser):

it's not C.

OpenStudy (jfraser):

set up the equation first, without numbers, then do the math one step at a time

OpenStudy (somy):

P1V1/T1=P2V2/T2 so if volume is constant then we can ignore it thus P1/T1=P2/V2 and so now solve it

OpenStudy (jfraser):

\[\frac{P_1}{T_1} = \frac{P_2}{T_2}\]so\[P_2 = \frac{P_1*T_2}{T_1}\]

OpenStudy (jfraser):

and convert the temp to kelvin, always

OpenStudy (somy):

btw if you take pressure as Kpa then the answer of the second pressure will be in Kpa

OpenStudy (anonymous):

i think i did it wrong.. i got (B)

OpenStudy (somy):

bingoo!

OpenStudy (jfraser):

it's B